Concentrated nitric acid reacts with solid copper to give nitrogen dioxide gas and dissolved copper ions according to the equation: Cu(s) + 4 H+(aq) + 2 NO3 - (aq) → 2 NO2(g) + Cu2+(aq) + 2 H2O(l) Suppose that 6.80 g of copper is consumed in this reaction and that the NO2 is collected at a pressure of 0.970 atm and a temperature of 45 oC. What volume of NO2 is produced?
Answer #1


Molar mass of Cu = 63.55 g/mol

mass of Cu = 6.8 g
mol of Cu = (mass)/(molar mass)
= 6.8/63.55
= 0.107 mol

According to balanced equation
mol of NO2 formed = (2/1)* moles of Cu
= (2/1)*0.107
= 0.214 mol

Given:
P = 0.97 atm
n = 0.214 mol
T = 45.0 oC
= (45.0+273) K
= 318 K

use:
P * V = n*R*T
0.97 atm * V = 0.214 mol* 0.08206 atm.L/mol.K * 318 K
V = 5.7571 L
Answer: 5.76 L